# Spirals and Beyond: Competitive Plane Search with Multi-Speed Agents

### v1

@@ Abstract
We consider the problem of minimizing the worst-case search time for a hidden point target in the plane using multiple mobile agents of differing speeds, all starting from a common origin. The search time is normalized by the target's distance to the origin, following the standard convention in competitive analysis. The objective is to minimize the maximum such normalized time over all possible target locations, known as the search cost. As a foundation for our main results, we extend the known result for a single unit-speed agent, which achieves a provably optimal search cost of approximately $\mathcal{U}_{1}=17.28935$ via a logarithmic spiral, to the setting of $n$ unit-speed agents. We design a symmetric spiral-based algorithm in which each agent follows a logarithmic spiral offset by equal angular phases. A key feature of this construction is that the resulting search cost is independent of which agent finds the target. We provide a closed-form upper bound $\mathcal{U}_{n}$ for the search cost in this setting, which forms the basis for our general result.
@@ 1 Introduction
We consider a related variant introduced in [49], where the target is a point in the plane. A point is exposed by an agent if it lies on the segment between the agent’s current and starting positions. The goal is to design a trajectory that minimises the worst-case target exposure time, normalized by the target's distance from the origin. The logarithmic spiral was first shown to be optimal among monotone periodic trajectories [49], and later proven unconditionally optimal in [65].
We extend the upper bound results to a team of multi-speed agents, a rarely studied generalization. First, we construct natural, non-intersecting logarithmic spiral trajectories where all agents incur equal cost, and identify speed regimes where agents do or do not contribute to exposing new points. As a corollary, we relate multi-speed performance to that of unit-speed agents, and give a condition, based on the geometric mean of the fastest agents’ speeds, for effective contribution. Second, motivated by the limitations of spiral trajectories in low-speed regimes, we introduce and analyze non-spiral algorithms and show that they can outperform the previously derived spirals in such settings. This suggests that spiral-search may not always be optimal, as has been shown or conjectured in other models.
@@ 1.1 Related Work
In two dimensions, a closely related problem is shoreline search, introduced by Bellman [17], where the target is an infinite line. The case with known shoreline distance to the agent's starting position was solved by Isbell [58], while the general “lost-in-a-forest” version (where the shape of the forest is part of the input) remains open [19, 46], with specific convex environments and general shapes having also been studied [54, 63]. For unknown distances, the best strategy known for shoreline search is a logarithmic spiral with competitive ratio $13.81$ [9, 47] and the best lower bounds known are $6.3972$ (unconditional) [8] and $12.5385$ (for cyclic trajectories) [66].
When the target is a point, which is the focus of this work, a logarithmic spiral achieves a search cost of $17.289$, and has been proven to be optimal. This result resolved a long-standing conjecture that spiral trajectories are best among all possible strategies. Spiral strategies in planar domains have been extensively studied [45, 47]. Gal [49] established optimality within the class of monotone, periodic strategies, and Langetepe [65] later proved unconditional optimality, confirming the conjecture. For known-distance variants, recent work provides exact and average-case optimal results [27].
Collaborative search with multiple agents has also received significant attention [3, 23, 33, 36, 38, 39, 57, 68]. The well-known ANTS problem [43] considers identical agents searching for an adversarially placed treasure in the plane, with performance depending on agent count and target distance. This extends earlier distance-cost models [10] and connects to other frameworks such as layered graph search [70] and ray-based exploration [7]. When two-dimensional search involves geometric specifications, the problem becomes more complex. Early work considered polygons [44] and disks [29], later extending to triangles, squares, and $\ell_{p}$ balls [11, 34, 52, 53]. Multi-agent extensions introduce coordination and fault-tolerance challenges [15, 20, 31], with some studies exploring worst- and average-case trade-offs [24].
@@ 2.1 Problem Definition, Notation & Terminology
For simplicity, we say that agent $i$ exposes a point $A\in\mathcal{D}$ if some point on $\tau_{i}$ exposes $A$. For each $A\in\mathcal{D}$, let $t_{A}$ be the earliest time some agent exposes $A$. The goal is to design trajectories that minimize the search cost, defined as $\sup_{A\in\mathcal{D}}\frac{t_{A}}{\left\lVert A\right\rVert},$ following the normalization of competitive analysis. Note that the origin is excluded from the search domain $\mathcal{D}$ so as to have a well-defined search cost.
@@
The notion of exposure corresponds to a well-known search/visibility criterion; a mobile agent sees a point if and only if the point lies between the agent's location and the origin. Our problem is to design trajectories for $n$ agents, considering their speed constraints, to locate a hidden target while minimizing the worst-case time required to expose a point relative to its distance from the origin.
Since trajectories are only relevant until a target is exposed, we assume each agent $i\in[n]$ moves at full speed $c_{i}$. Thus, we model agent trajectories as piecewise differentiable functions $\tau_{i}:\mathbb{R}\to\mathbb{R}^{2},$ where each $\tau_{i}$ satisfies $\lim_{t\rightarrow-\infty}\tau_{i}(t)=(0,0),$ with the understanding that $\tau_{i}(t)$ is a parametrization starting from the origin at $t=$-$\infty$ (that is, $t$ does not quantify time). These curves are assumed to be of infinite length to guarantee that every point in the plane is exposed by some agent. That is, for each $A\in\mathcal{D}$, there exist $i\in[n]$, $t_{A}\in\mathbb{R}$, and $\lambda\in[0,1]$ such that $A=\lambda\tau_{i}(t_{A}).$
The curve length from the origin to $\tau_{i}(t_{A})$ is
@@ 2.2 Main Contributions & Paper Organization
In this section, we summarize our main contributions, focusing on a simplified presentation of our results without explicitly describing the search algorithms. The primary focus of our work is to establish upper bounds to the search cost of $\mathcal{P}\text{-}\textsc{MSP}_{n}(c)$, i.e. the problem of searching the plane with arbitrary speed agents. As an essential component for the follow-up results, we first consider the simpler case $\mathcal{P}\text{-}\textsc{MSP}_{n}(\textbf{1})$, where all $n$ agents have unit speed. The proof of the following upper bound, based on a logarithmic-spiral type trajectory, generalizes the optimal solution for $\mathcal{P}\text{-}\textsc{MSP}_{1}(1)$ from [65]. The notation introduced in the following theorem is essential for presenting our general results.
Theorem 2.1.
For every $n\in\mathbb{Z}$, the search problem $\mathcal{P}\text{-}\textsc{MSP}_{n}(\textbf{1})$ admits a solution with search cost
@@
where $\kappa_{n}$ admits a closed-form expression (see Lemma 3.4).
@@
The proof of Theorem 2.2 appears in Section 4, and more specifically as the corollary of a sequence of lemmata, in Section 4.4. The proof constructs logarithmic-spiral trajectories tailored to the agents' speeds. Crucially, these trajectories introduce enough degrees of freedom to enable a competitive analysis that is agent-independent, a defining optimality condition for agents that contribute to search. Note that ordering the speeds in non-decreasing order simplifies the theorem's statement, as it maximizes the geometric mean over any subset of agents $[j]$, where $j\leq n$. Since $c_{\ell}\leq\mathcal{G}_{\ell}$ for all $\ell\in[n]$, Theorem 2.2 implies that an agent with speed $c_{\ell}$ contributes to exposing new targets only if $c_{\ell}$ is sufficiently large relative to the geometric mean of the speeds of all faster agents. This threshold is quantified by the factor
@@
These asymptotics will follow by the formula of $\mathcal{U}_{\ell}$ from Theorem 2.1 and Proposition 3.5 on page 3.5 regarding the behaviour of $\kappa_{n}$, with representative numerical values provided in Table 1. The following corollary follows immediately from Theorem 2.2.
@@
Corollary 2.3.
Searching for a point in the plane with $n$ agents of arbitrary speeds $c\in\mathbb{R}^{n}$ can be performed faster than searching with $k\leq n$ unit-speed agents, assuming that the geometric mean of the $n$ agents' speeds is at least $\mathcal{U}_{n}/\mathcal{U}_{k},$ where $\mathcal{U}_{i}$ is the search cost for $i$ unit-speed agents, as given in Theorem 2.1.
@@
$\frac{\sin\left(\phi\right)}{2\sin\left(\frac{\phi}{2}\right)}\cdot\mathcal{F}_{\phi},$
@@
$\gamma(\phi)=\frac{\sin\left(\phi\right)}{2\sin\left(\phi/2\right)}\frac{\mathcal{F}_{\phi}}{\mathcal{R}_{\phi}},$
where $\mathcal{F}_{\phi}$ and $\mathcal{R}_{\phi}$ are as defined in Theorems 2.4 and 2.5, respectively. Then, for every $\phi\in(0,\pi)$, the problem $\mathcal{P}\text{-}\textsc{MSP}_{2}(1,\gamma(\phi))$ admits a solution with search cost equal to $\mathcal{R}_{\phi}$, or also equal to
$\frac{\sin\left(\phi\right)}{2\sin\left(\phi/2\right)}\frac{\mathcal{F}_{\phi}}{\gamma(\phi)}.$
@@
We prove Theorem 2.7 in Section 7.
@@ 3 Searching the Plane with $n$ Unit-Speed Agents
Algorithm 1 is depicted in Figure 2. The execution of the algorithm begins at $t\to$-$\infty$. The condition $k>0$ is necessary for correctness, ensuring that every point in the plane is eventually exposed. The following lemma establishes that Algorithm 1 is indeed a feasible search strategy.
Lemma 3.1.
For every $k>0$ and $n\in\mathbb{Z}$, Algorithm 1 is feasible for $\mathcal{P}\text{-}\textsc{MSP}_{n}(\textbf{1})$.
Proof.
@@
$\left\lVert\,\mathrm{d}h(t)/\,\mathrm{d}t\right\rVert^{2}=e^{2(\alpha t+\beta)}\left(\left(\alpha\cos(t+\phi)-\sin(t+\phi)\right)^{2}+\left(\alpha\sin(t+\phi)+\cos(t+\phi)\right)^{2}\right)=e^{2(\alpha t+\beta)}\left(1+\alpha^{2}\right).$
@@
Define $t_{s}:=2\pi s$ for $s\in\mathbb{Z}$. That is, the target $A_{x}=(x,0)$ is first exposed when agent-0 reaches $f_{0}^{k}(t_{s})$ for some $s\in\mathbb{Z}$. It follows that $x\leq e^{kt_{s}}=e^{k(2\pi s)}$. Moreover, by Lemma 3.2, agent-0 reaches this point at time
@@
Minimizing the search cost expression in Lemma 3.3 with respect to $k=k(n)$ yields the following upper bounds for $\mathcal{P}\text{-}\textsc{MSP}_{n}(\textbf{1})$, along with the corresponding optimal expansion factors, denoted by $\kappa_{n}$. The next lemma provides a closed formula for $\kappa_{n}$.
Lemma 3.4.
For each $n\in\mathbb{N}$, the search cost of trajectories $\{f_{i}^{k}(t)\}_{i\in[n]}$ for $\mathcal{P}\text{-}\textsc{MSP}_{n}(\textbf{1})$, as seen in Lemma 3.3, is convex in expansion factor $k$. Moreover, the corresponding optimizers $k=\kappa_{n}$ are given as the unique positive root to $k^{3}+k=n/2\pi$, and therefore
@@
Proof of Theorem 2.1.
We use the trajectories from Algorithm 1 with expansion factor $\kappa_{n}$ from Lemma 3.4. By Lemma 3.3, the resulting search cost is as stated in the theorem. ∎
@@
Lastly, we show that $\mathcal{U}_{n}$ tends to $1$, as the number of agents $n$ grows. For that, we see from Lemma 3.4 that $\kappa_{n}=\Theta\left(n^{1/3}\right)$, and hence $\kappa_{n}/n=\Theta\left(n^{-2/3}\right)$. But then,

### v2

@@ Abstract
We consider the problem of minimizing the worst-case search time for a hidden point target in the plane using multiple mobile agents of differing speeds, all starting from a common origin. The search time is normalized by the target’s distance to the origin, following the standard convention in competitive analysis. The objective is to minimize the maximum such normalized time over all possible target locations, known as the search cost. As a foundation for our main results, we extend the known result for a single unit-speed agent, which achieves a provably optimal search cost of approximately $\mathcal{U}_{1}=17.28935$ via a logarithmic spiral, to the setting of $n$ unit-speed agents. We design a symmetric spiral-based algorithm in which each agent follows a logarithmic spiral offset by equal angular phases. A key feature of this construction is that the resulting search cost is independent of which agent finds the target. We provide a closed-form upper bound $\mathcal{U}_{n}$ for the search cost in this setting, which forms the basis for our general result.
@@ 1 Introduction
We consider a related variant introduced in [49], where the target is a point in the plane. A point is exposed by an agent if it lies on the line segment between the agent’s current position and the origin. The goal is to design a trajectory that minimises the worst-case target exposure time, normalized by the target’s distance from the origin. The logarithmic spiral was first shown to be optimal among monotone periodic trajectories [49], and later proven unconditionally optimal in [67].
We extend the upper bound results to a team of multi-speed agents, a rarely studied generalization. First, we construct natural, non-intersecting logarithmic spiral trajectories where all agents incur equal cost, and identify speed regimes where agents do or do not contribute to exposing new points. As a corollary, we relate multi-speed performance to that of unit-speed agents, and give a condition, based on the geometric mean of the fastest agents’ speeds, for effective contribution. Second, motivated by the limitations of spiral trajectories in low-speed regimes, we introduce and analyze non-spiral algorithms and show that they can outperform the previously derived spirals in such settings. This suggests that spiral-search may not always be optimal, as has been shown or conjectured in other models. Throughout the paper, statements of optimality are with respect to the underlying point search problem as defined above, the standard performance measure, and the agent speed regimes explicitly specified in each result.
@@ 1.1 Related Work
In two dimensions, a closely related problem is shoreline search, introduced by Bellman [17], where the target is an infinite line. The case with known shoreline distance to the agent’s starting position was solved by Isbell [60], while the general “lost-in-a-forest” version (where the shape of the forest is part of the input) remains open [19, 45], with specific convex environments and general shapes having also been studied [56, 65]. For unknown distances, the best strategy known for shoreline search is a logarithmic spiral with competitive ratio $13.81$ [8, 46] and the best lower bounds known are $6.3972$ (unconditional) [10] and $12.5385$ (for cyclic trajectories) [68].
When the target is a point, which is the focus of this work, a logarithmic spiral achieves a search cost of $17.289$, and has been proven to be optimal [67]. This result resolved a long-standing conjecture that spiral trajectories are best among all possible strategies. Spiral strategies in planar domains have been extensively studied [47, 46]. Gal [49] established optimality within the class of monotone, periodic strategies, and Langetepe [67] later proved unconditional optimality, confirming the conjecture. For known-distance variants, recent work provides exact and average-case optimal results [27]. Related geometric search models also admit non-spiral optimal trajectories; for example, recent work based on Fermat’s principle establishes optimal average-case strategies for disk inspection in a different setting [55].
Collaborative search with multiple agents has also received significant attention, see [2, 23, 33] and [36, 38, 39, 59, 70]. The well-known ANTS problem [43] considers identical agents searching for an adversarially placed treasure in the plane, with performance depending on agent count and target distance. This extends earlier distance-cost models [9] and connects to other frameworks such as layered graph search [72] and ray-based exploration [7]. When two-dimensional search involves geometric specifications, the problem becomes more complex. Early work considered polygons [44] and disks [29], later extending to triangles, squares, and $\ell_{p}$ balls [11, 34, 52, 54]. Multi-agent extensions introduce coordination and fault-tolerance challenges [15, 20, 31], with some studies exploring worst- and average-case trade-offs [24].
@@ 2.1 Problem Definition, Notation & Terminology
For simplicity, we say that agent $i$ exposes a point $A\in\mathcal{D}$ if some point on $\tau_{i}$ exposes $A$. For each $A\in\mathcal{D}$, let $t_{A}$ be the earliest time some agent exposes $A$. The goal is to design trajectories that minimize the search cost, defined as
$\sup_{A\in\mathcal{D}}\frac{t_{A}}{\left\lVert A\right\rVert},$
following the normalization of competitive analysis. [^1] The search cost corresponds to the notion of competitive ratio in competitive analysis. In some settings, the competitive ratio is defined with an additive term, or with an explicit restriction that excludes instances in which the reference cost becomes arbitrarily small (in our case, targets at very small distance $\left\lVert A\right\rVert$ from the origin), as a way to deal with pathological inputs. In geometric search, efficiency is instead measured directly by the ratio defining the search cost, as we do in this work. In our formulation, excluding the origin from the search domain is a sufficient and clean way to make this ratio well defined. This avoids both imposing an arbitrary positive lower bound on the target distance and introducing an additive term in the competitive ratio, which would be an unnecessary artefact in this model. Note that the origin is excluded from the search domain $\mathcal{D}$ so as to have a well-defined search cost. Intuitively, the denominator $\left\lVert A\right\rVert$ corresponds to the time required by an agent that knows the target location in advance to reach it, while the normalization yields a meaningful worst case measure by comparing the search time as a function of the distance of the hidden target from the origin, which would otherwise be unbounded.
@@
The notion of exposure corresponds to a well-known search/visibility criterion; a mobile agent sees a point if and only if the point lies between the agent’s location and the origin. Agents are modelled as points and do not interfere with one another; in particular, the exposure condition is defined per agent and is not affected by the locations of other agents (there is no notion of occlusion or blocking in the model). Our problem is to design trajectories for $n$ agents, considering their speed constraints, to locate a hidden target while minimizing the worst-case time required to expose a point relative to its distance from the origin.
Since trajectories are only relevant until a target is exposed, we assume each agent $i\in[n]$ moves at full speed $c_{i}$. Thus, we model agent trajectories as continuous, piecewise differentiable functions
$\tau_{i}:\mathbb{R}\to\mathbb{R}^{2},$
where each $\tau_{i}$ satisfies
$\lim_{t\to-\infty}\tau_{i}(t)=(0,0),$
with the understanding that $t$ is an abstract parameter and does not represent physical time. This convention should be read as a normalization, not as a physical requirement that the agent acts for an infinite amount of time. The parameter $t$ is a curve parameter, and physical time is obtained from arc length; saying that execution begins at $t\to$-$\infty$ simply means that the trajectory can be extended backwards so that, for every $\varepsilon>0$, the agent reaches distance at most $\varepsilon$ from the origin at some parameter value. In particular, this avoids introducing an arbitrary positive lower bound on the radius (or an arbitrary starting offset along the curve) when describing spiral-type trajectories.
We emphasize that we do not impose a common starting positive radius for the agents. For unit-speed trajectories, fixing a starting circle amounts only to a time shift and does not affect the competitive analysis, while introducing unnecessary normalization constants. More importantly, in the multi-speed setting such a requirement would be artificial: equal starting radii do not correspond to equal time or cost for agents of different speeds, and enforcing this constraint would obscure the geometric asymmetry that the algorithms in this paper are designed to exploit.
The piecewise differentiability assumption is standard in continuous motion models with bounded speed, as it ensures that instantaneous speed is well defined almost everywhere and that speed constraints can be meaningfully enforced. Curves $\tau_{i}$ are assumed to be of infinite length to guarantee that every point in the plane is exposed by some agent. That is, for each $A\in\mathcal{D}$, there exist $i\in[n]$, $t_{A}\in\mathbb{R}$, and $\lambda\in[0,1]$ such that $A=\lambda\tau_{i}(t_{A}).$
Distances are measured with respect to the standard Euclidean metric on $\mathbb{R}^{2}$, and speed bounds are interpreted with respect to arc length in this metric. Thus, the curve length from the origin to $\tau_{i}(t_{A})$ is
@@ 2.2 Main Contributions & Paper Organization
In this section, we summarize our main contributions, focusing on a simplified presentation of our results without explicitly describing the search algorithms. The primary focus of our work is to establish upper bounds to the search cost of $\mathcal{P}\text{-}\textsc{MSP}_{n}(c)$, i.e. the problem of searching the plane with arbitrary speed agents. As an essential component for the follow-up results, we first consider the simpler case $\mathcal{P}\text{-}\textsc{MSP}_{n}(\textbf{1})$, where all $n$ agents have unit speed. The proof of the following upper bound, based on a logarithmic-spiral type trajectory, generalizes the optimal solution for $\mathcal{P}\text{-}\textsc{MSP}_{1}(1)$ from [67]. The notation introduced in the following theorem, as well as the derived results, are essential for presenting our more general results.
Theorem 2.1
For every $n\in\mathbb{N}$, the search problem $\mathcal{P}\text{-}\textsc{MSP}_{n}(\textbf{1})$ admits a solution with search cost
@@
where $\kappa_{n}$ admits a closed-form expression, obtained as the root of an explicit cubic polynomial (see Lemma 4).
@@
The proof of Theorem 2.2 appears in Section 4, and more specifically as the corollary of a sequence of lemmata, in Section 4.4. The proof constructs logarithmic-spiral trajectories tailored to the agents’ speeds. Crucially, these trajectories introduce enough degrees of freedom to enable a competitive analysis that is agent-independent, a defining optimality condition for agents that contribute to search. Note that ordering the speeds in non-increasing order simplifies the theorem’s statement, as it maximizes the geometric mean over any subset of agents $[j]$, where $j\leq n$. Since $c_{\ell}\leq\mathcal{G}_{\ell}$ for all $\ell\in[n]$, Theorem 2.2 implies that an agent with speed $c_{\ell}$ contributes to exposing new targets only if $c_{\ell}$ is sufficiently large relative to the geometric mean of the speeds of all faster agents. This threshold is quantified by the factor
@@
These asymptotics will follow by the formula of $\mathcal{U}_{\ell}$ from Theorem 2.1 and Proposition 1 on page 1 regarding the behaviour of $\kappa_{n}$, with representative numerical values provided in Table 1. The following corollary follows immediately from Theorem 2.2.
@@
Corollary 1
Searching for a point in the plane with $n$ agents of arbitrary speeds $c\in\mathbb{R}^{n}$ can be performed faster than searching with $k\leq n$ unit-speed agents, assuming that the geometric mean of the $n$ agents’ speeds is at least $\mathcal{U}_{n}/\mathcal{U}_{k},$ where $\mathcal{U}_{i}$ is the search cost for $i$ unit-speed agents, as given in Theorem 2.1.
@@
$\cos\left(\phi/2\right)\cdot\mathcal{F}_{\phi},$
@@
$\gamma(\phi)=\cos\left(\phi/2\right)\frac{\mathcal{F}_{\phi}}{\mathcal{R}_{\phi}},$
where $\mathcal{F}_{\phi}$ and $\mathcal{R}_{\phi}$ are as defined in Theorems 2.3 and 2.4, respectively. Then, for every $\phi\in(0,\pi)$, the problem $\mathcal{P}\text{-}\textsc{MSP}_{2}(1,\gamma(\phi))$ admits a solution with search cost equal to $\mathcal{R}_{\phi}$, or also equal to
$\cos\left(\phi/2\right)\frac{\mathcal{F}_{\phi}}{\gamma(\phi)}.$
@@
The improvement achieved by the hybrid strategy occurs only in a narrow interval of speed ratios. For very small values of $c$, the slow agent cannot contribute meaningfully, and the bound coincides with the unit-speed spiral strategy; for larger values of $c$, the spiral-based bound $\mathcal{U}_{2}/\sqrt{c}$ dominates. The intermediate regime visible in Figure 1 arises from a numerical optimization of the cone angle in the hybrid construction, and reflects a geometric trade-off between restricting the slow agent to a narrow cone and increasing the cost of searching the complementary region. We prove Theorem 2.6 in Section 7.
@@ 3 Searching the Plane with $n$ Unit-Speed Agents
The logarithmic spiral guarantees that the distance of newly exposed points from the origin grows proportionally to the length of the trajectory. This property is necessary and sufficient for achieving a constant search cost. Using the same expansion factor across agents further guarantees that trajectories do not overlap.
Algorithm 1 is depicted in Figure 2. The execution of the algorithm begins at $t\to$-$\infty$. The condition $k>0$ is necessary for correctness, ensuring that every point in the plane is eventually exposed. For any fixed parameter value $t$, all agents have the same distance $e^{kt}$ from the origin, while their angular separation is exactly $2\pi/n$. Consequently, at every snapshot of the algorithm, all agents lie on a common circle and their positions form a regular $n$-gon (up to rotation and scaling as $t$ varies).
The following lemma establishes that Algorithm 1 is indeed a feasible search strategy.
Lemma 1
For every $k>0$ and $n\in\mathbb{N}$, Algorithm 1 is feasible for $\mathcal{P}\text{-}\textsc{MSP}_{n}(\textbf{1})$.
Proof
@@
$\left\lVert\,\mathrm{d}h(t)/\,\mathrm{d}t\right\rVert^{2}$ $=e^{2(\alpha t+\beta)}\left(\left(\alpha\cos(t+\phi)-\sin(t+\phi)\right)^{2}+\left(\alpha\sin(t+\phi)+\cos(t+\phi)\right)^{2}\right)$
$=e^{2(\alpha t+\beta)}\left(1+\alpha^{2}\right).$
@@
Define $t_{s}:=2\pi s$ for $s\in\mathbb{Z}$. That is, the target $A_{x}=(x,0)$ is first exposed when agent-0 reaches $f_{0}^{k}(t_{s})$ for some $s\in\mathbb{Z}$. It follows that $x\leq e^{kt_{s}}=e^{k(2\pi s)}$. Moreover, by Lemma 2, agent-0 reaches this point at time
@@
Minimizing the search cost expression in Lemma 3 with respect to $k=k(n)$ yields the following upper bounds for $\mathcal{P}\text{-}\textsc{MSP}_{n}(\textbf{1})$, along with the corresponding optimal expansion factors, denoted by $\kappa_{n}$. The next lemma provides a closed formula for $\kappa_{n}$.
Lemma 4
For each $n\in\mathbb{N}$, the search cost of trajectories $\{f_{i}^{k}(t)\}_{i\in[n]}$ for $\mathcal{P}\text{-}\textsc{MSP}_{n}(\textbf{1})$, as seen in Lemma 3, is convex in expansion factor $k$. Moreover, the corresponding optimizers $k=\kappa_{n}$ are given as the unique positive root to $k^{3}+k=n/2\pi$, and therefore
@@
Proof(of Theorem 2.1)
We use the trajectories from Algorithm 1 with expansion factor $\kappa_{n}$ from Lemma 4. By Lemma 3, the resulting search cost is as stated in the theorem. ∎
@@
Lastly, we show that $\mathcal{U}_{n}$ tends to $1$, as the number of agents $n$ grows. For that, we see from Lemma 4 that $\kappa_{n}=\Theta\left(n^{1/3}\right)$, and hence $\kappa_{n}/n=\Theta\left(n^{-2/3}\right)$. But then,